### v1

# A New Convergence Analysis of Plug-and-Play Proximal Gradient Descent Under Prior Mismatch

Abstract
In this work, we provide a new convergence theory for plug-and-play proximal gradient descent (PnP-PGD) under prior mismatch where the denoiser is trained on a different data distribution to the inference task at hand. To the best of our knowledge, this is the first convergence proof of PnP-PGD under prior mismatch. Compared with the existing theoretical results for PnP algorithms, our new results removed the need for several restrictive and unverifiable assumptions.
1 Introduction
Inverse problems involve the recovery of an unknown signal $\boldsymbol{x}\in\mathbb{R}^{n}$ from a set of noisy measurements $\boldsymbol{y}=\boldsymbol{A}\boldsymbol{x}+\boldsymbol{e}$, where $\boldsymbol{A}\in\mathbb{R}^{m\times n}$ is the measurement model and $\boldsymbol{e}$ is the noise. Inverse problems are often formulated and solved as optimization problems of form
$\boldsymbol{x}^{*}\in\underset{\boldsymbol{x}}{\arg\min}\lambda f(\boldsymbol{x})+g(\boldsymbol{x})$ (1)
where $f$ is a data-fidelity term, $g$ a regularization term and $\lambda>0$ a parameter that controls the strength of the regularization. Proximal algorithms are widely used for solving Eq. (1) when the regularizer is nonsmooth. For example, the PGD is a standard approach for solving Eq. (1) via the iterations
$\begin{cases}\boldsymbol{z}_{k+1}&=\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})\\ \boldsymbol{x}_{k+1}&=\mathsf{prox}_{\lambda g}(\boldsymbol{z}_{k+1}),\end{cases}$ (PGD)
The second step of (PGD) relies on the proximal operator defined as
$\mathsf{prox}_{\tau g}(\boldsymbol{z})\mathrel{\mathop{\ordinarycolon}}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{z}\|_{2}^{2}+\tau g(\boldsymbol{x})\right\},$ (2)
where $\tau>0$ controls the influence of $g$ and we fix $\tau=1$ in the following.
2 Theory
Work Method Target Denoiser? Assumptions
Convex functions
Schmidt et al. (2011) PGD $\epsilon$-optimal
Shoushtari et al. (2022) RED BD, BI
General Non-convex functions
Xu et al. (2020) PnP-PGD TD
Hurault et al. (2022b) PnP-PGD TD
Shoushtari et al. (2024) PnP-ADMM BD, BI
[…]
Table 1: Comparison of the convergence results in the PGD literature. Here TD = using Target Denoiser, BD = Bounded Denoiser, BI = Bounded Iterates.
This section presents a convergence analysis of PnP-PGD that accounts for the use of mismatched denoisers. It should be noted that the theoretical analysis of PGD and PnP-PGD has been previously discussed in (Schmidt et al., 2011; Xu et al., 2020; Hurault et al., 2022b; Tan et al., 2024). In addition,  Shoushtari et al. (2024) and Shoushtari et al. (2022) further theoretically learned about the prior mismatch in PnP-ADMM and RED with relatively restrictive and unverifiable assumptions of bounded-iterates (see Assumption 6 in (Shoushtari et al., 2024) and Assumption 4 in (Shoushtari et al., 2022)). The novelty of our work can be summarized as follows:
  • We analysis convergence and measure error bounds of the mismatched priors in PnP-PGD with relaxed assumptions;
  • Our theory accommodates nonconvex data fidelities, nonconvex regularizers and expansive denoisers.
2.1 PnP-PGD with Mismatched Denoiser
(PGD) alternates between a proximal operation on $g$ and a gradient descent step on $f$, when $f$ is differentiable. PnP methodology proposes to replace the proximal operator by a more general denosier $\mathsf{D}(\cdot)$, such as BM3D or CNN. Then, (PGD) can be summarized as
$\begin{cases}\boldsymbol{z}_{k+1}&=\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})\\ \boldsymbol{x}_{k+1}&=\mathsf{D}_{\sigma}(\boldsymbol{z}_{k+1}),\end{cases}$ (PnP-PGD)
where by analogy to $\tau$ in Eq. (2), we introduce the parameter $\sigma>0$ for controlling the relative strength of the denoiser $\mathsf{D}_{\sigma}$.
We denote the target distribution as $p_{\boldsymbol{x}}$, suppose that the denoiser $\mathsf{D}_{\sigma}$ is a MMSE estimator for the AWGN denoising problem
$\boldsymbol{v}=\boldsymbol{x}+\boldsymbol{e}\quad\text{with}\quad\boldsymbol{x}\sim p_{\boldsymbol{x}},\quad\boldsymbol{e}\sim\mathcal{N}(0,\sigma^{2}\boldsymbol{I}).$ (3)
The MMSE denoiser is the conditional mean estimator for Eq. (3) and can be expressed as
$\mathsf{D}_{\sigma}(\boldsymbol{v})\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[\boldsymbol{x}|\boldsymbol{v}]=\int_{\mathbb{R}^{n}}\boldsymbol{x}p_{\boldsymbol{x}|\boldsymbol{v}}(\boldsymbol{x}|\boldsymbol{v})\,\mathrm{d}\boldsymbol{x},$ (4)
where $p_{\boldsymbol{x|v}}(x|v)\propto G_{\sigma}(\boldsymbol{v}-\boldsymbol{x})p_{\boldsymbol{x}}(\boldsymbol{x})$, with $G_{\sigma}$ denoting the Gaussian density. Since the above integral is generally intractable, in practice, the denoiser corresponds to a deep model trained to minimize the mean squared error (MSE) loss
$\mathcal{L}(\mathsf{D}_{\sigma})=\mathbb{E}\left[\|\boldsymbol{x}-\mathsf{D}_{\sigma}(\boldsymbol{v})\|_{2}^{2}\right].$ (5)
where the solution of Eq. (5) satisfies Eq. (4).
When using a mismatched prior with a mismatched distribution $\widehat{p}(\boldsymbol{x})$, we replace the target denoiser $\mathsf{D}_{\sigma}$ in (PnP-PGD) with
$\boldsymbol{x}_{k+1}=\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z}_{k+1}),$
where $\widehat{\mathsf{D}}_{\sigma}$ is the mismatched MMSE denoiser.
2.2 Theoretical Analysis
Recently introduced Gradient-Step (GS) Denoiser (Hurault et al., 2022a; Cohen et al., 2021) writes
$\mathsf{D}_{\sigma}=\nabla h_{\sigma},$ (6)
with a potential
$h_{\sigma}\mathrel{\mathop{\ordinarycolon}}\boldsymbol{x}\to\frac{1}{2}\|\boldsymbol{x}\|^{2}-g_{\sigma}(\boldsymbol{x}).$ (7)
We first have that if $h_{\sigma}$ is convex, the GS denoiser $\mathsf{D}_{\sigma}=\nabla h_{\sigma}$ is linked to the proximal operator of some function (Gribonval and Nikolova, 2020)
$\phi_{\sigma}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}\mathrel{\mathop{\ordinarycolon}}\forall\boldsymbol{x}\in\mathbb{R}^{n},D_{\sigma}(\boldsymbol{x})\in\mathsf{prox}_{\phi_{\sigma}}(\boldsymbol{x})=\mathop{\arg\min}_{\boldsymbol{z}}\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{z}\|^{2}+\phi_{\sigma}(\boldsymbol{z}).$ (8)
The next proposition shows that, if the residual $\text{I}d-\mathsf{D}_{\sigma}$ is contractive, there exists a closed-form and smooth $\phi_{\sigma}$ such that $\mathsf{D}_{\sigma}=\mathsf{prox}_{\phi_{\sigma}}$ is single-valued.
Proposition 2.1 ((Hurault et al., 2022b))
Let $\mathcal{X}$ be an open convex subset of $\mathbb{R}^{n}$ and $g_{\sigma}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathbb{R}$ a $\mathcal{C}^{k+1}$ function with $k\geq 1$ and $\nabla g_{\sigma}$ $L$-Lipschitz with $L<1$. Then, for $h_{\sigma}\mathrel{\mathop{\ordinarycolon}}\boldsymbol{x}\to\frac{1}{2}\|\boldsymbol{x}\|^{2}-g_{\sigma}(\boldsymbol{x})$ and $\mathsf{D}_{\sigma}\mathrel{\mathop{\ordinarycolon}}=\nabla h_{\sigma}=\text{I}d-\nabla g_{\sigma}$,
  • (i)
    $h_{\sigma}$ is (1L)(1-L)-strongly convex and $\forall\boldsymbol{x}\in\mathcal{X}$, $J_{\mathsf{D}_{\sigma}}(\boldsymbol{x})$ is positive define;
  • (ii)
    $\mathsf{D}_{\sigma}$ is injective, $\mathrm{Im}(\mathsf{D}_{\sigma})$ is open and, $\forall\boldsymbol{x}\in\mathcal{X}$, $\mathsf{D}_{\sigma}(\boldsymbol{x})=\mathsf{prox}_{\phi_{\sigma}}(\boldsymbol{x})$, with $\phi_{\sigma}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ defined by
    $\phi_{\sigma}(\boldsymbol{x})\mathrel{\mathop{\ordinarycolon}}=\left\{\begin{array}[]{lc}\left.g_{\sigma}\left(\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})\right)\right)-\frac{1}{2}\left\|\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})-\boldsymbol{x}\right\|^{2}&\text{ if }\boldsymbol{x}\in\operatorname{Im}\left(\mathsf{D}_{\sigma}\right),\\ +\infty&\text{ otherwise },\end{array}\right$. (9)
  • (iii)
    $\phi_{\sigma}$ is $\mathcal{C}^{k}$ on $\mathrm{Im}(\mathsf{D}_{\sigma})$ and $\forall\boldsymbol{x}\in\mathsf{D}_{\sigma}$, $\nabla\phi_{\sigma}(\boldsymbol{x})=\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})-\boldsymbol{x}=\nabla g_{\sigma}(\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x}))$;
  • (v)
    $\nabla\phi_{\sigma}$ is $\frac{L}{1-L}$-Lipschitz on $\mathrm{Im}(\mathsf{D}_{\sigma})$.
Remark 2.1
Note that $\phi_{\sigma}$ being possibly nonconvex, and also note that $\mathsf{D}_{\sigma}$ is possibly not nonexpansive.
In the following, we study the convergence of the PnP-PGD with a target denoiser $\mathsf{D}_{\sigma}=\mathsf{prox}_{\phi_{\sigma}}$ that corresponds to the proximal operator of a nonconnvex regularization function $\phi_{\sigma}$, as well as a mismatched denoiser $\widehat{\mathsf{D}}_{\sigma}$ corresponds to a deep model trained to minimize the MSE loss. For that propose, we re-target the objective function
$F_{\lambda,\sigma}\mathrel{\mathop{\ordinarycolon}}=\lambda f+\phi_{\sigma},$ (10)
where $f$ is a data-fidelity term, $\lambda$ is a regularization parameter and $\phi_{\sigma}$ is defined as in Proposition 2.1 from the function $g_{\sigma}$ satisfying $\mathsf{D}_{\sigma}=\text{I}d-\nabla g_{\sigma}$. The analysis relies on the following set of assumptions that serve as sufficient conditions.
Assumption 2.1
$f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ differentiable with $L_{f}$-Lipschitz gradient.
This assumption is a standard assumption used in nonconvex optimization.
Assumption 2.2
$g_{\sigma}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathbb{R}$ a $\mathcal{C}^{2}$ function and $\nabla g_{\sigma}$ $L$-Lipschitz with $L<1$.
Instead of requiring $\mathsf{D}_{\sigma}$ is contractive, Assumption 2.2 assumes that $Id-\mathsf{D}_{\sigma}$ is contractive, which is a relaxed assumption as in (Ryu et al., 2019). Besides, If $\nabla g_{\sigma}$ has Lipschitz constant $L>1$, we can introduce $0<\alpha<\frac{1}{L}$ and relax the denoising operation by replacing $D_{\sigma}$ with $D_{\sigma}^{\alpha}=\alpha D_{\sigma}+(1-\alpha)Id=Id-\alpha\nabla g_{\sigma}$, then we can define $\phi_{\sigma}^{\alpha}$ from $\alpha g_{\sigma}$ and using the modified PnP-PGD $x_{k+1}=D_{\sigma}^{\alpha}(x_{k}-\lambda\nabla f(x_{k}))$, then we can get the same convergence result of scenario with $\nabla g_{\sigma}$ has Lipschitz constant $L<1$.
Assumption 2.3
$f$ and $g_{\sigma}$ bounded from below, then from Proposition 2.1, we get that $\phi_{\sigma}$ and $F_{\lambda,\sigma}$ are also bounded from below.
Assumption 2.3 implies that there exists $F_{\lambda,\sigma}^{*}>$-$\infty$ such that $F_{\lambda,\sigma}(\boldsymbol{x})\geq F_{\lambda,\sigma}^{*}$.
Assumption 2.4
The denoiser $\mathsf{D}_{\sigma}$ and $\widehat{\mathsf{D}}_{\sigma}$ have the same range $\mathsf{Im}(\mathsf{D}_{\sigma})$.
This assumption that the two image denoisers have the same range is also a relatively mild assumption. Ideally, both denoisers would have the same range corresponding to the set of desired images.
Assumption 2.5
The $k+1$-th output of mismatched denoiser $\boldsymbol{x}_{k+1}=\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z}_{k+1})$ is a $\epsilon_{k+1}$-optimal solution to the proximal problem in the sence of target denoiser $\mathsf{D}_{\sigma}$:
$\phi_{\sigma}(\boldsymbol{x}_{k+1})+\frac{1}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{z}_{k+1}\|^{2}\leq\min_{\boldsymbol{u}}\left(\phi_{\sigma}(\boldsymbol{u})+\frac{1}{2}\|\boldsymbol{u}-\boldsymbol{z}_{k+1}\|^{2}\right)+\epsilon_{k+1},\quad k=0,1,\ldots$ (11)
where $\boldsymbol{z}_{k+1}=\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})$.
By Proposition 2.1 we get that $H(\boldsymbol{x})=\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{z}\|^{2}+\phi_{\sigma}(\boldsymbol{x})$ is $\frac{1}{L+1}$-strongly convex (see Appendix for proof), which means
$H(\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z}_{k+1}))-H(\mathsf{D}_{\sigma}(\boldsymbol{z}_{k+1}))\geq\frac{1}{2(L+1)}\|\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z}_{k+1})-\mathsf{D}_{\sigma}(\boldsymbol{z}_{k+1})\|^{2}$ (12)
where $\mathsf{D}_{\sigma}(\boldsymbol{z}_{k+1})=\mathop{\arg\min}_{\boldsymbol{x}}H(\boldsymbol{x})$, then with Assumption 2.5 we can then get the equivalent form:
$\|\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z}_{k+1})-\mathsf{D}_{\sigma}(\boldsymbol{z}_{k+1})\|\leq\sqrt{2(L+1)H(\boldsymbol{x}_{k+1})-H(\boldsymbol{x}^{*})}\leq\sqrt{2(L+1)\epsilon_{k+1}},$ (13)
which is a common assumption as in Shoushtari et al. (2024, 2022).
We are now ready to present our convergence result under mismatched MMSE denoisers.
Theorem 2.1
Run PnP-PGD using a mismatched MMSE denoiser for $t\geq 1$ iterations under Assumption 2.1-2.5, with $\lambda L_{f}<1$, then we have
$\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right)+\frac{1}{t}\cdot\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\sum_{k=1}^{t}\epsilon_{k},$ (14)
where $\epsilon_{k}$ is the error term. In addition, if the sequence $\{\epsilon_{k}\}_{k\geq 1}$ is summable, we have that $\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|\to 0$ as $t\to\infty$.
Note that in Theorem 2.1, the usual condition on the stepsize becomes a condition on the regularization parameter $\lambda L_{f}<1$. The regularization trade-off parameter is then limited by the value of $L_{f}$. Even if this is not usual in optimization, we argue that this is a not a problem as the regularization strength is also regulated by the $\sigma$ parameter which we are free to tune manually.
Lemma 2.1
A sufficient condition to achieve the summable of $\{\epsilon_{k}\}_{k\geq 1}$ is that $\epsilon_{k}$ decreases as $O(1/k^{1+\delta})$ for any $\delta>0$.
When we replace the mismatched MMSE denoiser with the target MMSE denoiser, we recover the traditional PnP-PGD. To highlight the impact of the mismatch, we next provide the same statement but using the target denoiser.
Theorem 2.2
Run PnP-PGD using the MMSE denoiser for $t\geq 1$ iterations under Assumption 2.1-2.3, with $\lambda L_{f}<1$, then, we have
$\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{8}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right).$ (15)
3 Proofs
3.1 Warm-up
We prove first $\frac{1}{L+1}$-Strong-Convexity of $H(\boldsymbol{x})$
Proof According to Proposition 2.1: the denoiser $\mathsf{D}_{\sigma}=\text{I}d-\nabla g_{\sigma}$, where $\nabla g_{\sigma}$ is $L$-Lipschitz and $L<1$, and $\phi_{\sigma}$ is $\frac{L}{L+1}$ semi-convex, which means $\phi_{\sigma}(\boldsymbol{x})+\frac{1}{2}\cdot\frac{L}{L+1}\|\boldsymbol{x}\|^{2}$ is convex.
Then we can rewrite $H(\boldsymbol{x})$ with
$H(\boldsymbol{x})=\left[\phi_{\sigma}(\boldsymbol{x})+\frac{1}{2}\cdot\frac{L}{L+1}\|\boldsymbol{x}\|^{2}\right]+\left[\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{z}_{k+1}\|^{2}-\frac{1}{2}\cdot\frac{L}{L+1}\|\boldsymbol{x}\|^{2}\right],$ (16)
where the first term is convex, and remain quadratic term’s Hessian is $(1-\frac{L}{L+1})I=\frac{1}{L+1}I\succ 0$, thus the second term is also convex. In total, $H(\boldsymbol{x})$ is $\frac{1}{L+1}$ strongly convex, this complete the proof.
3.2 Proof of Lemma 2.1
Proof If there exists $C>0,\delta>0$, satisfy
$\epsilon_{k}\leq\frac{C}{k^{1+\delta}},$
then, since $1+\delta>1$, we have
$\sum_{k=1}^{\infty}\epsilon_{k}\leq C\sum_{k=1}^{\infty}\frac{1}{k^{1+\delta}}$
is convergence, which complete the proof.
3.3 Proof of Theorem 2.1
Before proof the main theorem, we first present some useful lemmas.
Lemma 3.1
Assume that Assumption 2.1-2.5 hold and let the sequence $\{\boldsymbol{x}_{k}\}$ be generated via iterations of PnP-PGD with the mismatched MMSE denoiser and $\lambda L_{f}<1$. Then for the objective function $F_{\lambda,\sigma}$ defined above, we have
$F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\leq F_{\lambda,\sigma}(\boldsymbol{x}_{k})+\epsilon_{k+1}-\frac{1-L_{f}}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}.$ (17)
Proof Under Assumption 2.5, denoiser $\widehat{\mathsf{D}}_{\sigma}$ is an approximation of target MMSE denoiser, at $k+1$-th iteration, suppose that $\boldsymbol{x}_{k+1}=\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z}_{k+1})$, we have
$\phi_{\sigma}(\boldsymbol{x}_{k+1})+\frac{1}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}\leq\min_{\boldsymbol{u}}\left(\phi_{\sigma}(\boldsymbol{u})+\frac{1}{2}\|\boldsymbol{u}-\boldsymbol{z}_{k+1}\|^{2}\right)+\epsilon_{k+1},$
Introducing the majorization-minimization (MM) function
$Q(\boldsymbol{x},\boldsymbol{y})=\lambda f(\boldsymbol{y})+\lambda\left<\boldsymbol{x}-\boldsymbol{y},\nabla f(\boldsymbol{y})\right>+\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{y}\|^{2}+\phi_{\sigma}(\boldsymbol{x}),$
satisfies
$Q(\boldsymbol{x}_{k+1},\boldsymbol{x}_{k})\leq\min_{\boldsymbol{x}}Q(\boldsymbol{x},\boldsymbol{x}_{k})+\epsilon_{k+1},$
since
$\min_{\boldsymbol{x}}Q(\boldsymbol{x},\boldsymbol{x}_{k})\leq Q(\boldsymbol{x}_{k},\boldsymbol{x}_{k})=F_{\lambda,\sigma}(\boldsymbol{x}_{k}),$
we have
$Q(\boldsymbol{x}_{k+1},\boldsymbol{x}_{k})\leq Q(\boldsymbol{x}_{k},\boldsymbol{x}_{k})+\epsilon_{k+1}.$
Then, note that $Q(\boldsymbol{x}_{k+1},\boldsymbol{x}_{k})\leq Q(\boldsymbol{x}_{k},\boldsymbol{x}_{k})+\epsilon_{k+1}$ implies that
$\lambda f(\boldsymbol{x}_{k})+\lambda\left<\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k},\nabla f(\boldsymbol{x}_{k})\right>+\frac{1}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}+\phi_{\sigma}(\boldsymbol{x}_{k+1})$ $\leq\lambda f(\boldsymbol{x}_{k})+\phi_{\sigma}(\boldsymbol{x}_{k})+\epsilon_{k+1}$ (18)
$\phi_{\sigma}(\boldsymbol{x}_{k+1})\leq\phi_{\sigma}(\boldsymbol{x}_{k})-\lambda\left<\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k},\nabla f(\boldsymbol{x}_{k})\right$> $-\frac{1}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}+\epsilon_{k+1}$ (19)
using the descent lemma
$f(\boldsymbol{x})\leq f(\boldsymbol{y})+\left<\boldsymbol{x}-\boldsymbol{y},\nabla f(\boldsymbol{y})\right>+\frac{1}{2t}\|\boldsymbol{x}-\boldsymbol{y}\|^{2}$
with $t=\frac{1}{L_{f}}$ and relation (19) we get
$F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})$ $=\lambda f(\boldsymbol{x}_{k+1})+\phi_{\sigma}(\boldsymbol{x}_{k+1})$
$\leq\phi_{\sigma}(\boldsymbol{x}_{k})-\lambda\left<\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k},\nabla f(\boldsymbol{x}_{k})\right>-\frac{1}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}+\epsilon_{k+1}+\lambda f(\boldsymbol{x}_{k})+$
$\qquad\lambda\left<\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k},\nabla f(\boldsymbol{x}_{k})\right>+\frac{L_{f}}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}$
$=F_{\lambda,\sigma}(\boldsymbol{x}_{k})+\epsilon_{k+1}-\frac{1-L_{f}}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2},$
this complete the proof.
Lemma 3.2
Assume that Assumption 2.1-2.5 holds, then for the $\epsilon$-optimal solution $\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z})=\hat{\boldsymbol{x}}$ to the proximal problem:
$H(\boldsymbol{u})=\mathop{\arg\min}_{\boldsymbol{u}}\frac{1}{2}\|\boldsymbol{u}-\boldsymbol{z}\|^{2}+\phi_{\sigma}(\boldsymbol{u}),$ (20)
we have
$\begin{cases}\nabla\phi_{\sigma}(\hat{\boldsymbol{x}})=\boldsymbol{z}-\hat{\boldsymbol{x}}-\boldsymbol{\delta}\\ \|\boldsymbol{\delta}\|\leq\sqrt{\frac{2\epsilon}{1-L}}.\end{cases}$ (21)
Proof For $\nabla^{2}H(\boldsymbol{x})=I+\nabla^{2}\phi_{\sigma}(\boldsymbol{x})$, with Proposition 2.1, we have $\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})=\boldsymbol{x}+\nabla\phi_{\sigma}(\boldsymbol{x})$. For mapping $\boldsymbol{y}=\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})$, the Jacobian is $J_{\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})}=I+\nabla^{2}\phi_{\sigma}(\boldsymbol{x})$, by inverse theorem we can get $J_{\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})}=\left[J_{\mathsf{D}_{\sigma}(\boldsymbol{y})}\right]^{-1}$, so
$\nabla^{2}H(\boldsymbol{x})=I+\nabla^{2}\phi_{\sigma}(\boldsymbol{x})=\left[J_{D_{\sigma}(\boldsymbol{y})}\right]^{-1}.$
Then, we know that $\mathsf{D}_{\sigma}=\nabla h_{\sigma}$, as a result $J_{\mathsf{D}_{\sigma}(\boldsymbol{y})}=\nabla^{2}h_{\sigma}(\boldsymbol{y})$.
Since $h_{\sigma}$ is (1L)(1-L)-strongly convex, which implies
$\nabla^{2}h_{\sigma}(\boldsymbol{y})\succeq(1-L)I,\quad(L<1)$
then we have
$\lambda_{\max}(\nabla^{2}H(\boldsymbol{x}))=\frac{1}{\lambda_{\min}J_{\mathsf{D}_{\sigma}(\boldsymbol{y})}}\leq\frac{1}{1-L}$
which means $H(\boldsymbol{x})$ is $\frac{1}{1-L}$-smooth. Since $H(\boldsymbol{x})$ is differentiable, denote that
[…]
then we get
$\nabla\phi_{\sigma}(\hat{\boldsymbol{x}})=\boldsymbol{z}-\hat{\boldsymbol{x}}-\boldsymbol{\delta},$
since $\hat{\boldsymbol{x}}$ is an $\epsilon$-optimal solution to $H(\boldsymbol{u})$, we have
$H(\hat{\boldsymbol{x}})-H(\boldsymbol{x}^{*})\leq\epsilon.$
Using the $\frac{1}{1-L}$-smoothness of $H(\boldsymbol{x})$, we have
$\|\nabla H(\hat{\boldsymbol{x}})\|^{2}\leq\frac{2}{1-L}\left(H(\hat{\boldsymbol{x}})-H(\boldsymbol{x}^{*})\right)\leq\frac{2\epsilon}{1-L},$
thus,
$\|\boldsymbol{\delta}\|^{2}\leq\frac{2\epsilon}{1-L}\quad\text{and},\quad\nabla\phi_{\sigma}(\hat{\boldsymbol{x}})=\boldsymbol{z}-\hat{\boldsymbol{x}}-\boldsymbol{\delta}.$
this complete proof.
Theorem Run PnP-PGD using a mismatched MMSE denoiser for $t\geq 1$ iterations under Assumption 2.1-2.5 with $\lambda L_{f}<1$, then we have
$\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right)+\frac{1}{t}\cdot\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\sum_{k=1}^{t}\epsilon_{k}.$ (22)
Proof At $k+1$-th iteration, for the output of mismatched denoiser $\boldsymbol{x}_{k+1}=\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z}_{k+1})$, by Lemma 3.2, we have
$\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})$ $=\lambda\nabla f(\boldsymbol{x}_{k+1})+\nabla\phi_{\sigma}(\boldsymbol{x}_{k+1})$
$=\lambda\nabla f(\boldsymbol{x}_{k+1})+\boldsymbol{z}_{k+1}-\boldsymbol{x}_{k+1}-\boldsymbol{\delta}_{k+1}$
$=\lambda\nabla f(\boldsymbol{x}_{k+1})+\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})-\boldsymbol{x}_{k+1}-\boldsymbol{\delta_{k+1}}$
$=(\boldsymbol{x}_{k}-\boldsymbol{x}_{k+1})+\lambda\left(\nabla f(\boldsymbol{x}_{k+1})-\nabla f(\boldsymbol{x}_{k})\right)-\boldsymbol{\delta_{k+1}},$
thus
$\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\leq\|\boldsymbol{x}_{k}-\boldsymbol{x}_{k+1}\|+\lambda\|\nabla f(\boldsymbol{x}_{k+1})-\nabla f(\boldsymbol{x}_{k})\|+\|\boldsymbol{\delta}_{k+1}\|,$
since $\nabla f$ is $L_{f}$-Lipschitz, we have
$\|\nabla F_{\sigma,\lambda}(\boldsymbol{x}_{k+1})\|\leq(1+\lambda L_{f})\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|+\|\boldsymbol{\delta}_{k+1}\|,$
using $(\mathbf{a}+\mathbf{b})^{2}\leq 2\mathbf{a}^{2}+2\mathbf{b}^{2}$ and Lemma 3.2, we have
$\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\|^{2}\leq 2(1+\lambda L_{f})^{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}+4\frac{\epsilon_{k+1}}{1-L},$
using Lemma 3.1, we have
$\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}\leq\frac{2}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{k})-F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\right)+\frac{2\epsilon_{k+1}}{1-L_{f}},$
combining these two relations,
$\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\|^{2}$ $\leq\frac{4(1+\lambda L_{f})^{2}}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{k})-F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\right)+\frac{4(1+\lambda L_{f})^{2}}{1-L_{f}}\epsilon_{k+1}+\frac{4\epsilon_{k+1}}{1-L}$
$\leq\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{k})-F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\right)+\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\epsilon_{k+1},$
summing over $k=0,1,\ldots,t-1$ gives
$\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right)+\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\sum_{k=1}^{t}\epsilon_{k}$
Then by averaging over $t\geq 1$ iterations, we have
$\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right)+\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\frac{1}{t}\cdot\sum_{k=1}^{t}\epsilon_{k},$
this complete the proof.
3.4 Proof of Theorem 2.2
Before the proof of main theorem, we first present a useful lemma.
Lemma 3.3
Assume that Assumption 2.1-2.3 hold and let the sequence $\{\boldsymbol{x}_{k}\}$ be generated via iterations of PnP-PGD using the MMSE denoiser with $\lambda L_{f}<1$. Then for the objective function $F_{\lambda,\sigma}$ defined above, we have
$F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\leq F_{\lambda,\sigma}(\boldsymbol{x}_{k})-\frac{1-L_{f}}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}.$ (23)
Proof Introducing the majorization-minimization (MM) function
$Q(\boldsymbol{x},\boldsymbol{y})=\lambda f(\boldsymbol{y})+\lambda\left<\boldsymbol{x}-\boldsymbol{y},\nabla f(\boldsymbol{y})\right>+\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{y}\|^{2}+\phi_{\sigma}(\boldsymbol{x}),$
we have
$Q(\boldsymbol{x},\boldsymbol{x})=F_{\lambda,\sigma}(\boldsymbol{x}),$
and
$\mathop{\arg\min}_{\boldsymbol{x}}Q(\boldsymbol{x},\boldsymbol{y})$ $=\mathop{\arg\min}_{\boldsymbol{x}}\lambda\left<\boldsymbol{x}-\boldsymbol{y},\nabla f(\boldsymbol{y})\right>+\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{y}\|^{2}+\phi_{\sigma}(\boldsymbol{x})$
$=\mathop{\arg\min}_{\boldsymbol{x}}\phi_{\sigma}(\boldsymbol{x})+\frac{1}{2}\|\boldsymbol{x}-(\boldsymbol{y}-\lambda\nabla f(\boldsymbol{y}))\|^{2}$
$=\mathrm{Prox}_{\phi_{\sigma}}\circ(Id-\lambda\nabla f)(\boldsymbol{y}),$
the $\mathop{\arg\min}$ is unique by Proposition 2.1 and by definition of the $\mathop{\arg\min}$, $\boldsymbol{x}_{k+1}=\mathrm{Prox}_{\phi_{\sigma}}\circ(Id-\lambda\nabla f)(\boldsymbol{x}_{k})$ implies that
$Q(\boldsymbol{x}_{k+1},\boldsymbol{x}_{k})\leq Q(\boldsymbol{x}_{k},\boldsymbol{x}_{k})=F_{\lambda,\sigma}(\boldsymbol{x}_{k}).$
Moreover, with $f$ being $L_{f}$-smooth, we have by the descent lemma, for any $t\leq\frac{1}{L_{f}}$ and any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{n}$,
$f(\boldsymbol{x})\leq f(\boldsymbol{y})+\left<\boldsymbol{x}-\boldsymbol{y},\nabla f(\boldsymbol{y})\right>+\frac{1}{2t}\|\boldsymbol{x}-\boldsymbol{y}\|^{2}.$
Note that $Q(\boldsymbol{x}_{k+1},\boldsymbol{x}_{k})\leq Q(\boldsymbol{x}_{k},\boldsymbol{x}_{k})$ implies that
$\phi_{\sigma}(\boldsymbol{x}_{k+1})\leq\phi_{\sigma}(\boldsymbol{x}_{k})-\lambda\left<\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k},\nabla f(\boldsymbol{x}_{k})\right>-\frac{1}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2},$
using above descent lemma with stepsize $t=\frac{1}{L_{f}}$, we get
$F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})$ $=\lambda f(\boldsymbol{x}_{k+1})+\phi_{\sigma}(\boldsymbol{x}_{k+1})$
$\leq\phi_{\sigma}(\boldsymbol{x}_{k})-\lambda\left<\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k},\nabla f(\boldsymbol{x}_{k})\right>-\frac{1}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}+\lambda f(\boldsymbol{x}_{k})+\lambda\left<\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k},\nabla f(\boldsymbol{x}_{k})\right$>
$\quad+\frac{L_{f}}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}$
$=F_{\lambda,\sigma}(\boldsymbol{x}_{k})-\frac{1-L_{f}}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2},$
this complete the proof.
Theorem Run PnP-PGD using the MMSE denoiser for $t\geq 1$ iterations under Assumption 2.1-2.3, with $\lambda L_{f}<1$, then, we have
$\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{8}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right).$ (24)
Proof By Proposition 2.1, for $\boldsymbol{x}\in\mathsf{Im}(\mathsf{D}_{\sigma})$, $\phi_{\sigma}$ is differentiable, and
$\nabla\phi_{\sigma}(\boldsymbol{x})=\mathsf{D}^{-1}_{\sigma}(\boldsymbol{x})-\boldsymbol{x},$
for the PnP-PGD iterates:
$\begin{cases}\boldsymbol{z}_{k+1}=\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})\\ \boldsymbol{x}_{k+1}=\mathsf{D}_{\sigma}(\boldsymbol{z}_{k+1})\end{cases}$
we first derivate the optimality condition:
$\nabla\phi_{\sigma}(\boldsymbol{x}_{k+1})$ $=\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x}_{k+1})-\boldsymbol{x}_{k+1}$
$=\boldsymbol{z}_{k+1}-\boldsymbol{x}_{k+1}$
$=\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})-\boldsymbol{x}_{k+1}$
and $\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})=\lambda\nabla f(\boldsymbol{x}_{k+1})+\nabla\phi_{\sigma}(\boldsymbol{x}_{k+1})$, we then have
$\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})$ $=\lambda\nabla f(\boldsymbol{x}_{k+1})+\nabla\phi_{\sigma}(\boldsymbol{x}_{k+1})$
$=\lambda\nabla f(\boldsymbol{x}_{k+1})+\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})-\boldsymbol{x}_{k+1}$
$=(\boldsymbol{x}_{k}-\boldsymbol{x}_{k+1})+\lambda\left(\nabla f(\boldsymbol{x}_{k+1})-\nabla f(\boldsymbol{x}_{k})\right)$
thus
$\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\|$ $\leq\|\boldsymbol{x}_{k}-\boldsymbol{x}_{k+1}\|+\lambda\|\nabla f(\boldsymbol{x}_{k+1})-\nabla f(\boldsymbol{x}_{k})\|$
$\leq\|\boldsymbol{x}_{k}-\boldsymbol{x}_{k+1}\|+\lambda L_{f}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|$
$\leq(1+\lambda L_{f})\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|$
then we have
$\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\|^{2}\leq(1+\lambda L_{f})^{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}.$
By Lemma 3.3, we have
$\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}\leq\frac{2}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{k})-F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\right),$
combining these two relations, we get
$\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\|^{2}$ $\leq\frac{2(1+\lambda L_{f})^{2}}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{k})-F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\right)$
$\leq\frac{8}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{k})-F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\right),$
summing over $k=1,2,\ldots,t$ gives
$\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{8}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right).$
Then by averaging over $t\geq 1$ iterations, we have
$\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{8}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right),$
this complete the proof.

### v2

@@
 Compared with the existing theoretical results for PnP algorithms, our new results removed the need for several restrictive and unverifiable assumptions.
+Moreover, we derive the convergence theory for equivariant PnP (EPnP) under the prior mismatch setting, proving that EPnP reduces error variance and explicitly tightens the convergence bound.
 1 Introduction
@@
 (1) via the iterations
-$\begin{cases}\boldsymbol{z}_{k+1}&=\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})\\ \boldsymbol{x}_{k+1}&=\mathsf{prox}_{\lambda g}(\boldsymbol{z}_{k+1}),\end{cases}$ (PGD)
+$\begin{cases}\boldsymbol{z}_{k+1}&=\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})\\ \boldsymbol{x}_{k+1}&=\mathsf{prox}_{\tau g}(\boldsymbol{z}_{k+1}),\end{cases}$ (PGD)
 The second step of (PGD) relies on the proximal operator defined as
@@
   • Our theory accommodates nonconvex data fidelities, nonconvex regularizers and expansive denoisers.
+  • We derive the convergence theory for equivariant PnP-PGD (EPnP) under the prior mismatch setting.
+We prove that EPnP reduces error variance and explicitly tightens the convergence bound for PnP-PGD.
 2.1 PnP-PGD with Mismatched Denoiser
@@
     $\mathsf{D}_{\sigma}$ is injective, $\mathrm{Im}(\mathsf{D}_{\sigma})$ is open and, $\forall\boldsymbol{x}\in\mathcal{X}$, $\mathsf{D}_{\sigma}(\boldsymbol{x})=\mathsf{prox}_{\phi_{\sigma}}(\boldsymbol{x})$, with $\phi_{\sigma}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ defined by
-    $\phi_{\sigma}(\boldsymbol{x})\mathrel{\mathop{\ordinarycolon}}=\left\{\begin{array}[]{lc}\left.g_{\sigma}\left(\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})\right)\right)-\frac{1}{2}\left\|\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})-\boldsymbol{x}\right\|^{2}&\text{ if }\boldsymbol{x}\in\operatorname{Im}\left(\mathsf{D}_{\sigma}\right),\\ +\infty&\text{ otherwise },\end{array}\right$.
+    $\phi_{\sigma}(\boldsymbol{x})\mathrel{\mathop{\ordinarycolon}}=\left\{\begin{array}[]{lc}\left(g_{\sigma}\left(\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})\right)\right)-\frac{1}{2}\left\|\mathsf{D}_{\sigma}^{-1}(\boldsymbol{x})-\boldsymbol{x}\right\|^{2}&\text{ if }\boldsymbol{x}\in\operatorname{Im}\left(\mathsf{D}_{\sigma}\right),\\ +\infty&\text{ otherwise },\end{array}\right$.
 (9)
@@
 where $\boldsymbol{z}_{k+1}=\boldsymbol{x}_{k}-\lambda\nabla f(\boldsymbol{x}_{k})$.
+Remark 2.2
 By Proposition 2.1 we get that $H(\boldsymbol{x})=\frac{1}{2}\|\boldsymbol{x}-\boldsymbol{z}\|^{2}+\phi_{\sigma}(\boldsymbol{x})$ is $\frac{1}{L+1}$-strongly convex (see Appendix for proof), which means
@@
 $\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(\boldsymbol{x}_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{8}{1-L_{f}}\left(F_{\lambda,\sigma}(\boldsymbol{x}_{0})-F_{\lambda,\sigma}^{*}\right).$ (15)
+2.3 Equivariant PnP-PGD with Mismatched Denoiser
+While prior mismatch is traditionally analyzed as a discrepancy in data distributions (e.g., training on faces vs.
+testing on landscapes) (Shoushtari et al., 2024), we argue that a fundamental source of mismatch stems from the inductive bias of the network architecture itself, which we term Structural Prior Mismatch.
+Seminal works have established that the architecture of a generator captures low-level image statistics, acting as an implicit prior (Ulyanov et al., 2018).
+However, standard CNN architectures impose rigid geometric constraints, such as fixed grid alignments, that do not strictly align with the continuous transformation groups of natural scenes (Cohen and Shashua, 2017).
+This misalignment manifests as spectral bias (Rahaman et al., 2019) and aliasing artifacts (Zhang, 2019), representing a systematic deviation $\mathcal{E}$ between the modeled prior $\hat{D}$ and the true prior $D^{*}$.
+Unlike stochastic noise, this structural error is often anisotropic (Antun et al., 2020), creating specific directions of instability in the PnP optimization landscape.
+Consequently, we employ the Equivariant PnP (EPnP) framework not merely as a heuristic, but as a principled geometric intervention to mitigate this structural mismatch.
+We demonstrate that EPnP provably reduces the mismatch error term $\epsilon$ defined in Theorem 2.1, providing a concrete mechanism to improve convergence reliability without retraining the prior.
+As shown in (Lenc and Vedaldi, 2015), natural images densities tend to be invariant to some set of transformations such as rotations or flips.
+To formalize these properties, we define the key notions of invariance and $\pi$-equivariance.
+Definition 2.1 (Invariance)
+We denote by $\mathrm{T}_{g}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\to\mathbb{R}^{n}$ a differentiable transformation of $\mathbb{R}^{n}$ and by $\mathcal{G}$ a measurable set of transformations of $\mathbb{R}^{n}$.
+A density $p$ on $\mathbb{R}^{n}$ is said to be invariant to a set of transformations $\mathcal{G}$ if $\forall g\in\mathcal{G}$, $p=p\circ\mathrm{T}_{g}$ a.e.
+Definition 2.2 ($\pi$-equivariance)
+We denote by $G\sim\pi$ a random variable of law $\pi$ on $\mathcal{G}$.
+A density $p$ on $\mathbb{R}^{n}$ is said to be $\pi$-equivariant if $\log p=\mathbb{E}_{G\sim\pi}\left[|\log(p\circ G)|\right]$ and $\mathbb{E}_{G\sim\pi}\left[|\log(p\circ G)|\right]<\infty$.
+Definition 2.2 relaxes the notion of invariance for a density in the following sense.
+If a density $p$ is invariant to each $g\in\mathcal{G}$, then $p$ is $\pi$-equivariant, whatever the distribution $\pi$ on $\mathcal{G}$.
+Assumption 2.6
+Suppose that the ideal density $p(x)$ on $\mathbb{R}^{n}$ is invariant to a set of transformations $\mathcal{G}$.
+All $\mathrm{T}_{g}\in\mathcal{G}$ are affine transformations with isometric linear parts, i.e.
+$\forall\mathrm{T}_{g}\in\mathcal{G}$, there exist $a\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\to\mathbb{R}^{n}$ a linear isometry and $c\in\mathbb{R}^{n}$, such that $\forall x\in\mathbb{R}^{n},\mathrm{T}_{g}(x)=ax+c$.
+Assumption 2.6 is verified for a set of transformations as detailed in Renaud et al. (2025).
+Assumption 2.7
+Suppose that the Gaussian noise $\boldsymbol{e}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ is isotropic.
+This assumption assumes the measurement noise follows an isotropic distribution (e.g., Gaussian white noise).
+This is a standard assumption in the PnP and RED literature (Shoushtari et al., 2024, 2022) and aligns with the physical properties of generic sensor noise.
+We note that for specific structured noise types (e.g., rain streaks or banding), this assumption may be relaxed, but such cases are outside the scope of this theoretical analysis.
+Proposition 2.2
+With Assumption 2.62.7, the target denoiser $\mathrm{D}^{*}_{\sigma}$ is $\pi$-equivairant, which means for $\forall\mathrm{T}_{g}\in\mathcal{G},x\in\mathbb{R}^{n}$:
+$\mathrm{T}_{g}^{-1}\circ\mathrm{D}^{*}_{\sigma}(\mathrm{T}_{g}\circ x)=\mathrm{D}^{*}_{\sigma}(x).$ (16)
+With Assumption 2.5 that the mismatched denoiser $\hat{\mathrm{D}}_{\sigma}(z)$ exists bias, here we denote as
+$\hat{\mathrm{D}}_{\sigma}(z)=\mathrm{D}^{*}_{\sigma}(z)+\mathcal{E}(z),$
+where $\|\mathcal{E}(z)\|^{2}\propto\epsilon$. We then denote the equivairant denoiser $\tilde{\mathrm{D}}_{\sigma}$ as
+$\tilde{\mathrm{D}}_{\sigma}(z)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{g\sim\pi}\left[\mathrm{T}_{g}^{T}\circ\hat{\mathrm{D}}_{\sigma}(\mathrm{T}_{g}\circ z)\right]=\int_{G}\mathrm{T}_{g}\cdot\hat{\mathrm{D}}_{\sigma}(\mathrm{T}_{g}z)d\pi(g),$
+where $\pi$ is the normalized Haar measure defined on group $G$.
+Before present our main result, we first present some definitions and a useful lemma as follows.
+Definition 2.3 (Isotropic Bias)
+The bias $\mathcal{E}$ satisfies $\mathcal{E}(\mathrm{T}_{g}x)=\mathrm{T}_{g}\mathcal{E}(x)$ for $\forall\mathrm{T}_{g}\in G$ a.e.
+Definition 2.4 (Anisotropic Bias)
+The vector value function $V_{x}(\mathrm{T}_{g})\mathrel{\mathop{\ordinarycolon}}=\mathrm{T}_{g}^{-1}\mathcal{E}(\mathrm{T}_{g}x)$ is non-constant for $\forall T_{g}\in G$ a.e.
+In other word, there exists a positive measurable set satisfying $V_{x}(\mathrm{T}_{g})\neq\mathrm{const}$.
+Lemma 2.2 (Equivariant Bias Reduction)
+For $\forall x\in\mathcal{X}$ and $g\sim G$, if the bias $\mathcal{E}(x)$ generated by PnP is anisotropic, then the bias $\tilde{\mathcal{E}}(x)$ generated by EPnP satisfies
+$\|\tilde{\mathcal{E}}(x)\|^{2}<\int_{G}\|\mathcal{E}(\mathrm{T}_{g}x)\|^{2}d\pi(g)=\mathbb{E}_{g}[\|\mathcal{E}(\mathrm{T}_{g}x)\|^{2}].$ (17)
+Remark 2.3
+We denote $\tilde{\epsilon}_{k}$ as the error of EPnP:
+$\|\tilde{\mathrm{D}}_{\sigma}(z_{k+1})-\mathrm{D}_{\sigma}(z_{k+1})\|\leq\sqrt{2(L+1)\tilde{\epsilon}_{k+1}},$ (18)
+and
+$\mathbb{E}_{g\sim\pi}\left[\|\hat{\mathrm{D}}_{\sigma}(\mathrm{T}_{g}z_{k+1})-\mathrm{D}_{\sigma}(\mathrm{T}_{g}z_{k+1})\|\right]\leq\sqrt{2(L+1)\hat{\epsilon}_{k+1}}.$ (19)
+the group averaged PnP error. Then we can conclude that
+$\tilde{\epsilon}_{k}=\hat{\epsilon}_{k}-\text{Var}_{g}<\hat{\epsilon}_{k}.$
+Now we present our main theoretical result.
+Theorem 2.3
+Run EPnP-PGD using a equivariant mismatched MMSE denoiser for $t\geq 1$ iterations under Assumption 2.1-2.5 with $\lambda L_{f}<1$, then we have
+$\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(x_{k})\|^{2}\leq\frac{1}{t}\sum_{k+1}^{t}\|\nabla F_{\lambda,\sigma}(x_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(x_{0})-F_{\lambda,\sigma}^{*}\right)+\frac{1}{t}\cdot\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\sum_{k=1}^{t}\tilde{\epsilon}_{k}.$ (20)
+Specifically, with Assumption 2.6 - 2.7, we have following convergence result for EPnP-PGD:
+$\sum_{k=1}^{t}\tilde{\epsilon}_{k}=\sum_{k=1}^{t}\left(\hat{\epsilon}_{k}-\underbrace{\text{Var}_{g,k}}_{\text{Gain}}\right)<\hat{\epsilon}_{k},$ (21)
+where $\text{Var}_{g,k}=\text{Var}_{g,k}(\mathrm{T}_{g}^{-1}\cdot\mathcal{E}(\mathrm{T}_{g}z_{k}))>0$, the more isotropic the bias becomes, the more robust the convergence trajectory.
 3 Proofs
@@
 Then for the objective function $F_{\lambda,\sigma}$ defined above, we have
-$F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\leq F_{\lambda,\sigma}(\boldsymbol{x}_{k})+\epsilon_{k+1}-\frac{1-L_{f}}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}.$ (17)
+$F_{\lambda,\sigma}(\boldsymbol{x}_{k+1})\leq F_{\lambda,\sigma}(\boldsymbol{x}_{k})-\frac{1-L_{f}}{2}\|\boldsymbol{x}_{k+1}-\boldsymbol{x}_{k}\|^{2}+\epsilon_{k+1}.$ (23)
 Proof Under Assumption 2.5, denoiser $\widehat{\mathsf{D}}_{\sigma}$ is an approximation of target MMSE denoiser, at $k+1$-th iteration, suppose that $\boldsymbol{x}_{k+1}=\widehat{\mathsf{D}}_{\sigma}(\boldsymbol{z}_{k+1})$, we have
@@
 this complete the proof.
+3.5 Proof of Proposition 2.2
+Proposition With Assumption 2.62.7, the target denoiser $\mathrm{D}^{*}_{\sigma}$ is $\pi$-equivairant, which means for $\forall\mathrm{T}_{g}\in\mathcal{G},x\in\mathbb{R}^{n}$:
+$\mathrm{T}_{g}^{-1}\circ\mathrm{D}^{*}_{\sigma}(\mathrm{T}_{g}\circ x)=\mathrm{D}^{*}_{\sigma}(x).$ (31)
+Proof We denote the clean image distribution as $p(x)$, noised image distribution as $p_{\sigma}(y)$:
+$p_{\sigma}(y)=\int p(x)\phi(y-x)dx,$
+for the transformed noised distribution:
+$p_{\sigma}(\mathrm{T}_{g}y)=\int p(x)\phi(\mathrm{T}_{g}y-x)dx,$
+let $x^{\prime}=\mathrm{T}_{g}^{T}x$, i.e.
+$x=\mathrm{T}_{g}x^{\prime}$.
+Since $\mathrm{T}_{g}$ is orthogonal, and $J|\det(\mathrm{T}_{g})|=1$, thus, $dx=dx^{\prime}$.
+Plug this into above equation, yields
+$p_{\sigma}(\mathrm{T}_{g}y)=\int p(\mathrm{T}_{g}x^{\prime})\phi(\mathrm{T}_{g}y-\mathrm{T}_{g}x^{\prime})dx^{\prime}.$
+since $p(\mathrm{T}_{g}x^{\prime})=p(x^{\prime})$ and $\phi(\mathrm{T}_{g}(y-x^{\prime}))=\phi(y-x^{\prime})$ from Assumption 2.62.7, we have
+$p_{\sigma}(\mathrm{T}_{g}y)=\int p(x^{\prime})\phi(y-x^{\prime})dx^{\prime}$
+which we can conclude that
+$p_{\sigma}(\mathrm{T}_{g}y)=p_{\sigma}(y),$
+which means the noised distribution is also invariant. Differentiating both sides w.r.t $y$ of above equation
+$\text{RHS}=\nabla_{y}p_{\sigma}(y)$
+$\text{LHS}=\nabla_{y}p_{\sigma}(\mathrm{T}_{g}y)=\mathrm{T}_{g}^{T}\nabla p_{\sigma}(\mathrm{T}_{g}y).$
+Then we can have
+$\nabla_{y}p_{\sigma}(y)=\mathrm{T}_{g}^{T}\nabla p_{\sigma}(\mathrm{T}_{g}y),$
+which is equivalent to
+$\mathrm{T}_{g}\nabla_{y}p_{\sigma}(y)$ $=\mathrm{T}_{g}\mathrm{T}_{g}^{T}\nabla p_{\sigma}(\mathrm{T}_{g}y)$
+$=\nabla p_{\sigma}(\mathrm{T}_{g}y).$
+Dividing both side by $p_{\sigma}(\mathrm{T}_{g}y)$:
+$\frac{\mathrm{T}_{g}\nabla_{y}p_{\sigma}(y)}{p_{\sigma}(y)}=\frac{\nabla p_{\sigma}(\mathrm{T}_{g}y)}{p_{\sigma}(\mathrm{T}_{g}y)}$
+which means
+$\nabla\log p_{\sigma}(\mathrm{T}_{g}y)=\mathrm{T}_{g}\nabla\log p_{\sigma}(y),$ (32)
+then by Tweedie formula, the target MMSE denoiser is defined as:
+$\mathrm{D}_{\sigma}^{*}(y)=y+\sigma^{2}\nabla\log p_{\sigma}(y)$
+and
+$\mathrm{D}_{\sigma}^{*}(\mathrm{T}_{g}y)=\mathrm{T}_{g}y+\sigma^{2}\nabla\log p_{\sigma}(\mathrm{T}_{g}y).$
+Using Eq. (32) we have
+$\mathrm{D}_{\sigma}^{*}(\mathrm{T}_{g}y)=\mathrm{T}_{g}y+\sigma^{2}\mathrm{T}_{g}\nabla\log p_{\sigma}(y)$
+$\mathrm{D}_{\sigma}^{*}(\mathrm{T}_{g}y)=\mathrm{T}_{g}\left(y+\sigma^{2}\nabla\log p_{\sigma}(y)\right)=\mathrm{T}_{g}\mathrm{D}_{\sigma}^{*}(y),$
+this complete the proof.
+3.6 Proof of Lemma 2.2
+Lemma (Equivariant Bias Reduction) For $\forall x\in\mathcal{X}$ and $g\sim G$, if the bias $\mathcal{E}(x)$ generated by PnP is anisotropic, then the bias $\tilde{\mathcal{E}}(x)$ generated by EPnP satisfies
+$\|\tilde{\mathcal{E}}(x)\|^{2}<\int_{G}\|\mathcal{E}(\mathrm{T}_{g}x)\|^{2}d\pi(g)=\mathbb{E}_{g}[\|\mathcal{E}(\mathrm{T}_{g}x)\|^{2}].$ (33)
+Proof First we have
+$\tilde{\mathcal{E}}(x)$ $=\int_{G}\mathrm{T}_{g}^{T}\cdot\hat{\mathrm{D}}_{\sigma}(\mathrm{T}_{g}x)d\pi(g)-\mathrm{D}_{\sigma}^{*}(x)$
+$=\int_{G}\mathrm{T}_{g}^{T}\cdot\left[\mathrm{D}_{\sigma}^{*}(\mathrm{T}_{g}x)+\mathcal{E}(\mathrm{T}_{g}x)\right]d\pi(g)-\mathrm{D}_{\sigma}^{*}(x)$
+$=\int_{G}\left[\mathrm{T}_{g}^{T}\cdot\mathrm{D}_{\sigma}^{*}(\mathrm{T}_{g}x)\right]d\pi(g)+\int_{G}\left[\mathrm{T}_{g}^{T}\cdot\mathcal{E}(\mathrm{T}_{g}x)\right]d\pi(g)-\mathrm{D}_{\sigma}^{*}(x),$
+from Proposition 2.2, we have that the target denoiser $\mathrm{D}_{\sigma}^{*}$ is equivariant, i.e., $\mathrm{T}_{g}^{T}\cdot\mathrm{D}_{\sigma}^{*}(\mathrm{T}_{g}\cdot x)=\mathrm{D}_{\sigma}^{*}(x)$, then we can conclude that
+$\tilde{\mathcal{E}}(x)=\mathbb{E}_{g\sim\pi}[]T_{g}^{-1}\mathcal{E}(T_{g}x)],$
+using variance decomposition we have
+$\|\tilde{\mathcal{E}}(x)\|^{2}=\mathbb{E}_{g}[\|\mathrm{T}_{g}^{-1}\cdot\mathcal{E}(\mathrm{T}_{g}x)\|^{2}]-\text{Var}_{g}(\mathrm{T}_{g}^{-1}\cdot\mathcal{E}(\mathrm{T}_{g}x))$
+since $\mathrm{T}_{g}$ is isometric transform, i.e. $\|\mathrm{T}_{g}^{-1}\cdot v\|=\|v\|$, thus above equation is equivalent to
+$\|\tilde{\mathcal{E}}(x)\|^{2}=\mathbb{E}_{g}[\|\mathcal{E}(\mathrm{T}_{g}x)\|^{2}]-\text{Var}_{g}(\mathrm{T}_{g}^{-1}\cdot\mathcal{E}(\mathrm{T}_{g}x)),$
+since bias $\mathcal{E}(x)$ is anisotropic, we have
+$\text{Anisotropy}\implies\text{Var}_{g}(\mathrm{T}_{g}^{-1}\cdot\mathcal{E}(\mathrm{T}_{g}x))>0,$
+thus
+$\|\tilde{\mathcal{E}}(x)\|^{2}=\mathbb{E}_{g}[\|\mathcal{E}(\mathrm{T}_{g}x)\|^{2}]-\underbrace{\text{Var}(...)}_{>0}<\mathbb{E}_{g}[\|\mathcal{E}(\mathrm{T}_{g}x)\|^{2}]$
+this complete the proof.
+3.7 Proof of Theorem 2.3
+the proof is similar to Theorem 2.1. At $k+1$-th iteration, for the output of mismatched denoiser $x_{k+1}=\tilde{\mathrm{D}}_{\sigma}(z_{k+1})$, by Lemma 3.2, we have
+$\nabla F_{\lambda,\sigma}(x_{k+1})$ $=\lambda\nabla f(x_{k+1})+\nabla\tilde{\phi}_{\sigma}(x_{k+1})$
+$=\lambda\nabla f(x_{k+1})+z_{k+1}-x_{k+1}-\tilde{\delta}_{k+1}$
+$=\lambda\nabla f(x_{k+1})+x_{k}-\lambda\nabla f(x_{k})-x_{k+1}-\tilde{\delta}_{k+1}$
+$=(x_{k}-x_{k+1})+\lambda\left(\nabla f(x_{k+1})-\nabla f(x_{k})\right)-\tilde{\delta}_{k+1},$
+thus
+$\nabla F_{\lambda,\sigma}(x_{k+1})\leq\|x_{k}-x_{k+1}\|+\lambda\|\nabla f(x_{k+1})-\nabla f(x_{k})\|+\|\tilde{\delta}_{k+1}\|,$
+since $\nabla f$ is $L_{f}$-Lipschitz, we have
+$\|\nabla F_{\sigma,\lambda}(x_{k+1})\|\leq(1+\lambda L_{f})\|x_{k+1}-x_{k}\|+\|\tilde{\delta}_{k+1}\|,$
+using $(\mathbf{a}+\mathbf{b})^{2}\leq 2\mathbf{a}^{2}+2\mathbf{b}^{2}$ and Lemma 3.2, we have
+$\|\nabla F_{\lambda,\sigma}(x_{k+1})\|^{2}\leq 2(1+\lambda L_{f})^{2}\|x_{k+1}-x_{k}\|^{2}+4\frac{\tilde{\epsilon}_{k+1}}{1-L},$
+using Lemma 3.1, we have
+$\|x_{k+1}-x_{k}\|^{2}\leq\frac{2}{1-L_{f}}\left(F_{\lambda,\sigma}(x_{k})-F_{\lambda,\sigma}(x_{k+1})\right)+\frac{2\tilde{\epsilon}_{k+1}}{1-L_{f}},$
+combining these two relations,
+$\|\nabla F_{\lambda,\sigma}(x_{k+1})\|^{2}$ $\leq\frac{4(1+\lambda L_{f})^{2}}{1-L_{f}}\left(F_{\lambda,\sigma}(x_{k})-F_{\lambda,\sigma}(x_{k+1})\right)+\frac{4(1+\lambda L_{f})^{2}}{1-L_{f}}\tilde{\epsilon}_{k+1}+\frac{4\tilde{\epsilon}_{k+1}}{1-L}$
+$\leq\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(x_{k})-F_{\lambda,\sigma}(x_{k+1})\right)+\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\tilde{\epsilon}_{k+1},$
+summing over $k=0,1,\ldots,t-1$ gives
+$\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(x_{k})\|^{2}\leq\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(x_{0})-F_{\lambda,\sigma}^{*}\right)+\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\sum_{k=1}^{t}\tilde{\epsilon}_{k}.$
+Then by averaging over $t\geq 1$ iterations, we have
+$\min_{1\leq k\leq t}\|\nabla F_{\lambda,\sigma}(x_{k})\|^{2}\leq\frac{1}{t}\sum_{k=1}^{t}\|\nabla F_{\lambda,\sigma}(x_{k})\|^{2}\leq\frac{1}{t}\cdot\frac{16}{1-L_{f}}\left(F_{\lambda,\sigma}(x_{0})-F_{\lambda,\sigma}^{*}\right)+\left(\frac{16}{1-L_{f}}+\frac{4}{1-L}\right)\frac{1}{t}\cdot\sum_{k=1}^{t}\tilde{\epsilon}_{k},$
+then by Lemma 2.2, we have
+$\tilde{\epsilon}_{k}=\hat{\epsilon}_{k}-\text{Var}_{g}<\hat{\epsilon}_{k},$
+which complete the proof.